OpenStudy (anonymous):

Please help me evaluate! tan(sin^-1(-5/9))

OpenStudy (anonymous):

Trigonometric Functions

OpenStudy (ranga):



OpenStudy (anonymous):

what would i do to find tan(x)??

OpenStudy (anonymous):

is there a formula?

OpenStudy (anonymous):

To find AB use pythagoras

OpenStudy (anonymous):

but how can i find anything without A and B having any numbers to enter into a formula?

OpenStudy (ranga):

Let sin^-1(-5/9) = x sin(x) = -5/9 tan(x) = opposite / adjacent = AC / AB = 5 / AB. AB can be found from AB^2 + AC^2 = AC^2

OpenStudy (anonymous):

Or tan(x) = sin(x)/cos(x) But u still have to find adjacent for it

OpenStudy (anonymous):

I meant Pythagoras theorem u can't use PYTHOGARAS because he is already dead I know it was a lame joke

OpenStudy (anonymous):

lol okay so i now know that i need to do ab^2 + ac^2 = ac^2 but how do i get the numbers? :(

OpenStudy (ranga):

AC = 5, BC = 9 (see the diagram).

OpenStudy (anonymous):

In a Right-angled triangle , Sum of the squares of two sides = the square of the hyptenuse

OpenStudy (ranga):

It should be AB^2 + AC^2 = BC^2

OpenStudy (adamaero):

SOH CAH TOA

OpenStudy (anonymous):

I got 10.3

OpenStudy (anonymous):

You wrote the formula incorrectly

OpenStudy (anonymous):

am i wrong :O

OpenStudy (anonymous):

i meant ali

OpenStudy (anonymous):

I will have to check

OpenStudy (anonymous):

okay because if i am right I'm still not sure what to do after lol

OpenStudy (anonymous):

Are u screwed with your basics

OpenStudy (ranga):

AB^2 = 9^2 - 5^2. AB = ?

OpenStudy (anonymous):

https://www.khanacademy.org/

OpenStudy (anonymous):

well I'm doing an online class with no textbook or teacher so what i do know i either taught myself or learned from youtube and math isn't my best subject, so i guess so

OpenStudy (anonymous):

no problem but don't jump the gun then u won't understand future concepts if you are not knowing the fundamentals

OpenStudy (anonymous):

ranga, i got 2sqrt14

OpenStudy (ranga):

Find AB. Find tan(x). Stick a negative sign in front of it because inverse sine of -5/9 is a negative angle and is in the fourth quadrant where tan is negative.

OpenStudy (anonymous):

oh wow, i see what happened there!

OpenStudy (ranga):

Yes, AB = 2sqrt(14). tan(x) = AC/AB = - 5 / (2sqrt(14))

OpenStudy (ranga):

Alright.

OpenStudy (anonymous):

would the answer be -5sqrt14/28? because i have to multiply by sqrt 14

OpenStudy (ranga):

Yes. Most teachers and textbooks don't like a radical in the denominator and so you will have to rationalize the denominator by multiplying top and bottom by sqrt(14). You will get -5sqrt(14) / 28.

OpenStudy (anonymous):

awesome, thank you so much!! :)

OpenStudy (ranga):

You are welcome.